# CogSteer: Cognition-Inspired Selective Layer Intervention for Efficiently Steering Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) achieve remarkable performance through pretraining on extensive data. This enables efficient adaptation to diverse downstream tasks. However, the lack of interpretability in their underlying mechanisms limits the ability to effectively steer LLMs for specific applications. In this work, we investigate the intrinsic mechanisms of LLMs from a cognitive perspective using eye movement measures. Specifically, we analyze the layer-wise correlation between human cognitive indicators and LLM representations. Building on these insights, we propose a heuristic approach for selecting the optimal steering layer to modulate LLM semantics. To this end, we introduce an efficient selective layer intervention based on prominent parameter-efficient fine-tuning methods, which conventionally adjust either all layers or only the final layer. Additionally, we present an implicit layer contrastive intervention during inference to steer LLMs away from toxic outputs. Extensive experiments on natural language understanding, reasoning, and generation tasks, conducted on GPT-2, LLaMa2-7B, and Mixtral-7B, demonstrate the effectiveness and efficiency of our approach. As a model-agnostic framework, it enhances the interpretability of LLMs while improving efficiency for safe deployment.
@@ 1 Introduction
Large Language Models (LLMs) Dubey et al. (2024); Yang et al. (2024); Guo et al. (2025) have demonstrated strong capabilities in natural language understanding and reasoning Wei et al. (2022); Zhao et al. (2023) through pretraining on large datasets, followed by instruction tuning and alignment with human values Wei et al. (2021); Ouyang et al. (2022). Consequently, LLMs achieve excellent performance on downstream tasks with fine-tuning. However, their lack of interpretability and transparency limits the development of efficient fine-tuning and inference methods.
@@
In this work, we introduce a novel interpretability analysis method that leverages eye movement data Luke and Christianson (2018); Hollenstein et al. (2020); Colman et al. (2022) collected by cognitive researchers to study human reading behaviors. Through correlation experiments, we find that LLM hidden states exhibit a strong correlation with human gaze, peaking in the middle layers. Using eye movement measures such as fixation and regression Rayner (1998) as human-interpretable indicators, we observe a hierarchical progression in LLMs, from initial syntactic and semantic processing to deeper integration and final prediction. Additionally, a comparison of correlation results between natural reading and task-specific reading suggests that the upper layers of LLMs are more capable of reasoning. Furthermore, advanced LLMs such as LLaMa, which incorporate instruction tuning and reinforcement learning from human feedback (RLHF), demonstrate enhanced reasoning capabilities, even in the middle layers.
@@
To achieve this, we refine prominent PEFT methods, which traditionally adjust either all layers or only the last layer. As shown in Figure 1, our proposed CogSteer framework first identifies the most suitable layer $M$ for semantic intervention based on the task. Instead of fine-tuning all layers or only the last layer, our method fine-tunes only the selected layer, enabling LLMs to better adapt to specific tasks and datasets. The number of learnable parameters in CogSteer is significantly reduced, requiring only $1/N$ of the parameters in LLMs, thereby improving parameter efficiency. Furthermore, we propose an implicit layer contrastive intervention method during inference, which efficiently identifies and steers semantics toward safer generation directions to evaluate the effectiveness of our proposed selective layer intervention.
@@
(1) We are the first to propose leveraging eye movement measures to analyze LLM behavior. We publicly release the probing code to facilitate further research on interpretability from a cognitive perspective. (2) Through correlation analysis, we demonstrate a hierarchical progression in LLMs and introduce a heuristic steering layer selection method for efficient layer intervention. (3) Extensive experiments validate the effectiveness of our proposed method across various language understanding, reasoning, and generation tasks, contributing to the development of efficient and explainable foundation models.
@@ 2 Related Work
research is essential for uncovering the mechanisms of LLMs and ensuring their safe and trustworthy deployment as foundation models. Various studies have analyzed how knowledge is stored in LLMs Goldowsky-Dill et al. (2023); Stolfo et al. (2023); Rai et al. (2024); Bills et al. (2023); Geva et al. (2022), focusing on concepts such as knowledge neurons Dai et al. (2022) and circuits Conmy et al. (2023); Yao et al. (2024). Moreover, Wang et al. (2024a) reveal that GPT-2 predicts tokens more similarly to humans than shallow language models but lacks engagement with LLMs. Unlike previous works, our method investigates the interpretability of LLMs through human-interpretable indicators based on eye movement theory, enabling a more precise understanding and control of model behavior (see § 3 and  4.1).
@@ 3 From Human Gaze to LLM Behavior
Correlation Analysis. Let $S_{j}$ denote the $j$-th sentence, consisting of $n_{j}$ words $w_{1},w_{2},\dots,w_{n_{j}}$. For each word $w_{i}$ in sentence $S_{j}$, we consider five eye movement measures: $e_{i}^{(k)},k\in\{\mathrm{\textit{std}},\,\mathrm{\textit{ffd}},\,\mathrm{\textit{gd}},\,\mathrm{\textit{trt}},\,\mathrm{\textit{gpt}}\}$, where each measure represents a scalar value. The hidden state at layer $l$ of the LLM for word $w_{i}$ is denoted as $\mathbf{h}_{l,i}\in\mathbb{R}^{d}$, where $d$ is the dimensionality of the hidden states.
@@
Results and Finding 1. Based on the correlation calculation in Eq. 1, we first analyze the correlation results for the natural reading task. To facilitate interpretation, we divide the layers of both models into three equal groups: premature, middle, and mature. Figures 2 (upper) and  7 present the correlation results between various eye movement measures and the hidden states of the LLMs, illustrating how these values evolve across layers.
@@
Finding 1. (Layer-wise functionality) LLM hidden states exhibit a strong correlation with human gaze, peaking in the middle bucket and again in the mature bucket. This pattern suggests a hierarchical progression from initial syntactic and semantic processing to deeper integration and final prediction.
@@ 4.1 Heuristic Steering Layer Selection
From a task-oriented perspective, we apply PEFT methods and inference-only methods to the candidate layers in the middle bucket, using a small portion of data, like the validation set, to search for and select the best-performing layer that suits the task scenario. Formally, given $M$ as the best layer for intervention, $J$ represents a set of candidate layers in the middle bucket $\left(\frac{N}{3}\leq M\leq\frac{2N}{3}\right)$. $D$ denotes the validation set. We search for the layer $M$ that yields the best task score or loss performance, as follows:
$M=\arg\max_{j\in J}S\left(D;P\left(\cdot\mid x_{t}\right)\right).$ (2)
@@ 4.3 Layer Intervention during Inference
where $v_{c}^{M}$ and $v_{o}^{M}$ are the contextualized value vectors of the contrast LLM and the original LLM at the best layer $M$ for semantic intervention (see Elhage et al. (2021) for mathematical derivation). We then update the value vector in the layer $M$ of the original LLM:
@@ 5.1 Datasets and Evaluation
Evaluation For the evaluation on the GLUE benchmark, we report both validation-set and test-set F1 scores for QQP and MRPC, while accuracy is used for all other tasks. Additionally, we use the RealToxicityPrompts (RTP) dataset Gehman et al. (2020). Following prior work Li et al. (2023); Leong et al. (2023), we sample 2,122 toxic prompts. For each toxic prompt in the RTP dataset, we generate 25 continuations and evaluate their toxicity using the Perspective API, which assigns a toxicity score to each continuation, with higher scores indicating a greater likelihood of toxicity. Finally, we use the average maximum toxicity as our evaluation metric. Specifically, we compare the toxicity scores and detoxification margins obtained by applying our methods to different layers of the models.
@@ 5.2 Models and Baselines
We evaluate our efficient semantic steering methods using three sizes of GPT-2 models, as discussed in § 3, along with LLaMa2-7B. We select GPT-2 and LLaMa2-7B because the former represents earlier classical LLMs, while the latter exemplifies modern LLMs. To ensure generalizability, we also evaluate the Mixtral-7B Jiang et al. (2024) model to assess performance across tasks. Since applying semantic intervention to either the last layer or all layers of LLMs is a conventional approach, we use these two methods as baselines for comparison. The implementation details are in Appendix A.
@@ 5.3 Evaluation on GLUE Benchmark
We first present the performance of our proposed selected layer intervention method on the GLUE Benchmark in Table 1. It can be observed that by selecting the optimal layer to steer semantics in LLMs for a specific task, all three LLMs achieve comparable or even superior results compared to conventional all-layer intervention while introducing only $1/N$ of the parameters ($N$ is the number of layers in the LLMs). Specifically, LLaMa2-7B achieves an absolute increase of +1.8 on average in the test set, while Mistral-7B achieves an absolute increase of +4.7 on average in the validation set and +5.9 in the test set. Moreover, GPT2 achieves comparable or better results on 5 out of 8 tasks in GLUE, LLaMa2 on 4 tasks, and Mixtral on 7 tasks.
Additionally, we find that the optimal layer for our proposed method consistently falls within the middle bucket across all LLMs. For GPT2-L, the best-performing layer is L19 for most tasks, while for LLaMA2-7B and Mixtral-7B, it is L14 and L12, respectively. This aligns with our interpretability analysis and findings based on eye movement measures (see § 3). Furthermore, the best layer remains consistent across both the validation and test sets, demonstrating the effectiveness of the heuristic steering layer selection approach (§ 4.1).
Lastly, from a task perspective, we find that paraphrase identification (MRPC, QQP) and natural language inference (MNLI, QNLI, RTE, WNLI) achieve the highest average improvement of +2.0 across all LLMs compared to full-layer fine-tuning (e.g., +28.7 for Mixtral-7B and +16.5 for LLaMA2-7B on the RTE test set). This suggests that, given the complex structure of LLMs, fine-tuning all layers for semantic steering does not always yield the best downstream task performance and can result in parameter redundancy, where certain layers become less active in making accurate predictions. Additionally, the conventional approach of fine-tuning only the last layer of an LLM, based on its proximity to the prediction output, is not optimal. Instead, in practice, the intervention layer should first be identified to better steer semantics and improve prediction accuracy for a specific task.
@@ 5.4 Analysis on Language Toxification
We evaluate our selective layer intervention via fine-tuning using the language toxification task. Figure 3 (green bars in the upper chart) shows the toxicity score for inserting a toxic adapter into each layer of the LLMs, with the last layer and full layers (red line) used as baselines. The best layer $M$ in the middle bucket for semantic intervention is $L23$ for GPT-2 and $L13$ for LLaMa2-7B, competing with the results from both the last layer (+4.5% for GPT & +23% for LLaMa) and full layers (+0.9% for GPT & +2.8% for LLaMa). This indicates that the conventional intervention is not optimal, demonstrating the effectiveness of our proposed selective layer intervention approach, which saves considerable computational resources and time for fine-tuning.
Furthermore, layer intervention in the mature bucket, including the last layer (used as a baseline), is not remarkable when compared to full-layer intervention for both GPT-2 and LLaMa2-7B models. According to our behavioral analysis in Section 3, the layers in the mature bucket focus more on reasoning and factual knowledge, involving less token processing, which makes them less suitable for semantic steering. Interestingly, for the GPT-2 model, we observe that layers $L1$-$L4$ and $L6$-$L8$ in the premature bucket also outperform the full-layer intervention baseline, whereas this phenomenon is not observed in LLaMa2-7B. We hypothesize that this difference may be due to the distinct adapter training schemes. For GPT-2, we use the vanilla Adapter as a plug-in, which directly affects token distributions and allows successive minor changes as tokens pass through layers, following the early exit theory Elbayad et al. (2020); Schuster et al. (2022). However, for LLaMa2, we use the LLaMa-Adapter, which directly modifies attention. Edits to the attention modules gradually affect token distributions Geva et al. (2021); Elhage et al. (2021).
@@ 5.5 Analysis on Language Detoxification
For Mistral-7B, the optimal layer for semantic intervention in the middle bucket is L15 for the detoxification task. Although the detoxification margin scores for L23 and L32 in the mature bucket are slightly higher than that of L15 (+0.14 on average), the scores in the middle bucket are more stable. We hypothesize that the differences among GPT-2, LLaMa2, and Mixtral-7B in the mature bucket, where Mixtral-7B achieves better detoxification performance, stem from variations in post-training. As an advanced LLM incorporating additional human value alignment for safety, Mixtral-7B benefits more from interventions in the mature bucket, leading to improved detoxification performance.

### v3

@@ Abstract
Large Language Models (LLMs) achieve remarkable performance through pretraining on extensive data. This enables efficient adaptation to diverse downstream tasks. However, the lack of interpretability in their underlying mechanisms limits the ability to effectively steer LLMs for specific applications. In this work, we investigate the intrinsic mechanisms of LLMs from a cognitive perspective using eye movement measures. Specifically, we analyze the layer-wise correlation between human cognitive indicators and LLM representations. Building on these insights, we propose a heuristic approach for selecting the optimal steering layer to modulate LLM semantics. To this end, we introduce an efficient selective layer intervention based on prominent parameter-efficient fine-tuning methods, which conventionally adjust either all layers or only the final layer. Additionally, we present an implicit layer contrastive intervention during inference to steer LLMs away from toxic outputs. Extensive experiments on natural language understanding, reasoning, and generation tasks, conducted on GPT-2, Llama2-7B, and Mistral-7B, demonstrate the effectiveness and efficiency of our approach. As a model-agnostic framework, it enhances the interpretability of LLMs while improving efficiency for safe deployment.
@@ 1 Introduction
Large Language Models (LLMs) Dubey et al. (2024); Yang et al. (2024); Guo et al. (2025) have demonstrated strong capabilities in natural language understanding and reasoning Zhao et al. (2024, 2023); Wei et al. (2022); Yin et al. (2025); Zeng et al. (2025) through pretraining on large datasets, followed by instruction tuning and alignment with human values Wei et al. (2021); Ouyang et al. (2022). Consequently, LLMs achieve excellent performance on downstream tasks with fine-tuning. However, their lack of interpretability and transparency limits the development of efficient fine-tuning and inference methods.
@@
In this work, we introduce a novel interpretability analysis method that leverages eye movement data Luke and Christianson (2018); Hollenstein et al. (2020); Colman et al. (2022) collected by cognition researchers to study human reading behaviors. Through correlation experiments, we find that LLM hidden states exhibit a strong correlation with human gaze, peaking in the middle layers. Using eye movement measures such as fixation and regression Rayner (1998) as human-interpretable indicators, we observe a hierarchical progression in LLMs, from initial syntactic and semantic processing to deeper integration and final prediction. Additionally, a comparison of correlation results between natural reading and task-specific reading suggests that the upper layers of LLMs are more capable of reasoning. Furthermore, advanced LLMs such as Llama, which incorporate instruction tuning and reinforcement learning from human feedback (RLHF), demonstrate enhanced reasoning capabilities, even in the middle layers.
@@
To achieve this, we refine prominent Parameter-Efficient Fine-Tuning (PEFT) methods, which traditionally adjust either all layers or only the last layer. As shown in Figure 1, our proposed CogSteer framework first identifies the most suitable layer $M$ for semantic intervention based on the task. Instead of fine-tuning all layers or only the last layer, our method fine-tunes only the selected layer, enabling LLMs to better adapt to specific tasks and datasets. The number of learnable parameters in CogSteer is significantly reduced, requiring only $1/N$ of the parameters in LLMs, thereby improving parameter efficiency. Furthermore, we propose an implicit layer contrastive intervention method during inference, which efficiently identifies and steers semantics toward safer generation directions to evaluate the effectiveness of our proposed selective layer intervention.
@@
(1) We are the first to propose leveraging eye movement measures to analyze the layer-wise behavior of LLMs. We publicly release the probing code  [^1] https://github.com/Ethanscuter/CogSteer to facilitate further research on interpretability from a cognitive perspective. (2) Through correlation analysis, we demonstrate a hierarchical progression in LLMs and introduce a heuristic steering layer selection method for efficient layer intervention. (3) Extensive experiments validate the effectiveness of our proposed method across various language understanding, reasoning, and generation tasks, contributing to the development of efficient and explainable foundation models.
@@ 2 Related Work
research is essential for uncovering the mechanisms of LLMs and ensuring their safe and trustworthy deployment as foundation models. Various studies have analyzed how knowledge is stored in LLMs Goldowsky-Dill et al. (2023); Stolfo et al. (2023); Rai et al. (2024); Bills et al. (2023); Geva et al. (2022), focusing on concepts such as knowledge neurons Dai et al. (2022) and circuits Conmy et al. (2023); Yao et al. (2024). Moreover, Wang et al. (2024a) reveal that GPT-2 predicts tokens more similarly to humans than shallow language models but lacks engagement with LLMs. Oh and Schuler (2023) analyzes how model scale and training data influence surprisal-based predictions of reading time. Gao et al. (2023) investigates the alignment between model and human attention. Unlike previous works, our method investigates the interpretability of LLMs through human-interpretable indicators based on eye movement theory, enabling a more precise understanding and control of model behavior (see § 3 and  4.1).
@@ 3 From Human Gaze to LLM Behavior
Correlation Analysis. Let $S_{j}$ denote the $j$-th sentence, consisting of $n_{j}$ words $w_{1},w_{2},\dots,w_{n_{j}}$. For each word $w_{i}$ in sentence $S_{j}$, we consider five eye movement measures: $e_{i}^{(k)},k\in\{\mathrm{\textit{sfd}},\,\mathrm{\textit{ffd}},\,\mathrm{\textit{gd}},\,\mathrm{\textit{trt}},\,\mathrm{\textit{gpt}}\}$, where each measure represents a scalar value. The hidden state at layer $l$ of the LLM for word $w_{i}$ is denoted as $\mathbf{h}_{l,i}\in\mathbb{R}^{d}$, where $d$ is the dimensionality of the hidden states.
@@
Results and Finding 1. Based on the correlation calculation in Eq. 1, we first analyze the correlation results for the natural reading task. To facilitate interpretation, we divide the layers of both models into three equal groups: premature, middle, and mature. The tripartite division of layers into premature, middle, and mature buckets is motivated by recent interpretability studies Geva et al. (2021); Schuster et al. (2022); Chuang et al. (2023); Zhang et al. (2024), which consistently observe a coarse-grained functional structure in LLMs: lower layers predominantly capture syntactic patterns, middle layers encode semantic and contextual information, and upper layers are often involved in reasoning and factual retrieval. Figures 2 (upper) and  7 present the correlation results between various eye movement measures and the hidden states of the LLMs, illustrating how these values evolve across layers.
@@
Finding 1. (Layer-wise functionality) LLM hidden states exhibit a strong correlation with human gaze, characterized by three distinct rises across layers. This pattern suggests a hierarchical progression from initial syntactic and semantic processing to deeper integration and final prediction.
@@ 4.1 Heuristic Steering Layer Selection
From a task-oriented perspective, we apply PEFT methods and inference-only methods to the candidate layers in the middle bucket, using a small portion of data, like the validation set, to search for and select the best-performing layer that suits the task scenario. Formally, given $M$ as the best layer for intervention, $J$ represents a set of candidate layers in the middle bucket $\left(\frac{N}{3}\leq M\leq\frac{2N}{3}\right)$. $D$ denotes the validation set. We search for the layer $M^{\prime}$ that yields the best task score or loss performance, as follows:
$M^{\prime}=\arg\max_{l\in M}Score\left(D;P\left(\cdot\mid x_{t},l\right)\right).$ (2)
@@ 4.3 Layer Intervention during Inference
where $v_{c}^{M}$ and $v_{o}^{M}$ are the contextualized value vectors of the contrast LLM and the original LLM at the best layer $M$ for semantic intervention. The contextualized value vectors Elhage et al. (2021) are derived as follows:
$A^{\ell,h}$ $=\varphi\left(\frac{\left(\mathbf{x}^{\ell-1}W_{Q}^{\ell,h}\right),\left(\mathbf{x}^{\ell-1}W_{K}^{\ell,h}\right)^{T}}{\sqrt{d/H}}+M^{\ell,h}\right),$ (5)
$\mathbf{a}^{\ell}$ $=\sum_{h=1}^{H}A^{\ell,h}\left(\mathbf{x}^{\ell-1}W_{V}^{\ell,h}\right)W_{O}^{\ell,h}=\sum_{h=1}^{H}\mathbf{v}^{\ell,h}W_{O}^{\ell,h}.$ (6)
Specifically, $\left(\mathbf{x}^{\ell-1}W_{V}^{\ell,h}\right)$ represents the attention-weighted, context-sensitive value vector for head $h$. $\mathbf{v}_{i}^{\ell,h}\in\mathbb{R}^{d}$ is the contextualized value vector at position $i$. We then update the value vector in the layer $M$ of the original LLM:
@@ 5.1 Datasets and Evaluation
Evaluation For the evaluation on the GLUE benchmark, we report both validation-set and test-set F1 scores for QQP and MRPC, while accuracy is used for all other tasks. Additionally, we use the RealToxicityPrompts (RTP) dataset Gehman et al. (2020). Following prior work Li et al. (2023); Leong et al. (2023), we sample 2,122 toxic prompts. For each toxic prompt in the RTP dataset, we generate 25 continuations and evaluate their toxicity using the Perspective API  [^3] https://perspectiveapi.com/, which assigns a toxicity score to each continuation, with higher scores indicating a greater likelihood of toxicity. Finally, we use the average maximum toxicity as our evaluation metric. Specifically, we compare the toxicity scores and detoxification margins obtained by applying our methods to different layers of the models.
@@ 5.2 Models and Baselines
We evaluate our efficient semantic steering methods using three sizes of GPT-2 models, as discussed in § 3, along with Llama2-7B. We select GPT-2 and Llama2-7B because the former represents earlier classical LLMs, while the latter exemplifies modern LLMs. To ensure generalizability, we also evaluate the Mistral-7B Jiang et al. (2023) model to assess performance across tasks. Since applying semantic intervention to either the last layer or all layers of LLMs is a conventional approach, we use these two methods as baselines for comparison. The implementation details are in Appendix A.
@@ 5.3 Evaluation on GLUE Benchmark
We first present the performance of our proposed selected layer intervention method on the GLUE Benchmark in Table 1. It can be observed that by selecting the optimal layer to steer semantics in LLMs for a specific task, all three LLMs achieve comparable or even superior results compared to conventional all-layer intervention while introducing only $1/N$ of the parameters ($N$ is the number of layers in the LLMs). Specifically, Llama2-7B achieves an absolute increase of +1.8 on average in the test set, while Mistral-7B achieves an absolute increase of +4.7 on average in the validation set and +5.9 in the test set. Moreover, GPT2 achieves comparable or better results on 5 out of 8 tasks in GLUE, Llama2 on 4 tasks, and Mistral on 7 tasks.
Additionally, we find that the optimal layer for our proposed method consistently falls within the middle bucket across all LLMs. For GPT2-L, the best-performing layer is L19 for most tasks, while for Llama2-7B and Mistral-7B, it is L14 and L12, respectively. This aligns with our interpretability analysis and findings based on eye movement measures (see § 3). Furthermore, the best layer remains consistent across both the validation and test sets, demonstrating the effectiveness of the heuristic steering layer selection approach (§ 4.1).
Lastly, from a task perspective, we find that paraphrase identification (MRPC, QQP) and natural language inference (MNLI, QNLI, RTE, WNLI) achieve the highest average improvement of +2.0 across all LLMs compared to full-layer fine-tuning (e.g., +28.7 for Mistral-7B and +16.5 for Llama2-7B on the RTE test set). This suggests that, given the complex structure of LLMs, fine-tuning all layers for semantic steering does not always yield the best downstream task performance and can result in parameter redundancy, where certain layers become less active in making accurate predictions. Additionally, the conventional approach of fine-tuning only the last layer of an LLM, based on its proximity to the prediction output, is not optimal. Instead, in practice, the intervention layer should first be identified to better steer semantics and improve prediction accuracy for a specific task.
@@ 5.4 Analysis on Language Toxification
We evaluate our selective layer intervention via fine-tuning using the language toxification task. Figure 3 (green bars in the upper chart) shows the toxicity score for inserting a toxic adapter into each layer of the LLMs, with the last layer and full layers (red line) used as baselines. The best layer $M$ in the middle bucket for semantic intervention is $L23$ for GPT-2 and $L13$ for Llama2-7B, competing with the results from both the last layer (+4.5% for GPT-2 & +23% for Llama2) and full layers (+0.9% for GPT-2 & +2.8% for Llama2). This indicates that the conventional intervention is not optimal, demonstrating the effectiveness of our proposed selective layer intervention approach, which saves considerable computational resources and time for fine-tuning.
Furthermore, layer intervention in the mature bucket, including the last layer (used as a baseline), is not remarkable when compared to full-layer intervention for both GPT-2 and Llama2-7B models. According to our behavioral analysis in Section 3, the layers in the mature bucket focus more on reasoning and factual knowledge, involving less token processing, which makes them less suitable for semantic steering. Interestingly, for the GPT-2 model, we observe that layers $L1$-$L4$ and $L6$-$L8$ in the premature bucket also outperform the full-layer intervention baseline, whereas this phenomenon is not observed in Llama2-7B. We hypothesize that this difference may be due to the distinct adapter training schemes. For GPT-2, we use the vanilla Adapter as a plug-in, which directly affects token distributions and allows successive minor changes as tokens pass through layers, following the early exit theory Elbayad et al. (2020); Schuster et al. (2022). However, for Llama2, we use the LLaMa-Adapter Zhang et al. (2023), which directly modifies attention. Edits to the attention modules gradually affect token distributions Geva et al. (2021); Elhage et al. (2021).
@@ 5.5 Analysis on Language Detoxification
For Mistral-7B, the optimal layer for semantic intervention in the middle bucket is L15 for the detoxification task. Although the detoxification margin scores for L23 and L32 in the mature bucket are slightly higher than that of L15 (+0.14 on average), the scores in the middle bucket are more stable. We hypothesize that the differences among GPT-2, Llama2, and Mistral-7B in the mature bucket, where Mistral-7B achieves better detoxification performance, stem from variations in post-training. As an advanced LLM incorporating additional human value alignment for safety, Mistral-7B benefits more from interventions in the mature bucket, leading to improved detoxification performance.